# Monochromatic products in random integer sets

### v1

@@ 2.1. Lack of forbidden configurations
For $n\in{\mathbb{N}}$, a function $p=p(n)\leq n^{-1/9}\omega(n)^{-1}$, and $S\sim[n]_{p}$ we have that a.a.s. there is a forbidden configuration in $S$.
@@
For $n\in{\mathbb{N}}$, $p\leq n^{-1/9}\omega(n)^{-1}$, and $S\sim[n]_{p}$ we have that a.a.s. there is a forbidden configuration in $S\cap(n^{1/9},n].$
@@
Let $e_{1},\dots,e_{k}\in\mathbb{N}$ with $1\leq e_{1}\leq e_{2}\leq\dots\leq e_{k}$ and $1\leq t\leq n$. Let $e\coloneq\sum_{i=1}^{k}e_{i}$. Then the number of $k$-tuples $(a_{1},\dots,a_{k})$ such that $t\leq a_{i}\leq n$ for all $2\leq i\leq k$ such that $t\leq a_{i}\leq n$ for all $1\leq i\leq k$ with $e_{i}\geq 2$, and $\prod_{i=1}^{k}a_{i}^{e_{i}}\leq n$ is at most
@@
As $\prod_{i=1}^{k}a_{i}^{e_{i}}\leq n$, and $a_{i}\geq t$ for each $i\in[k]$, any valid choice of $(a_{1},\ldots,a_{k})$ will have $\ell:=\prod_{i=1}^{k}a_{i}\leq nt^{k-e}$. After choosing $\ell$ (at most $nt^{k-e}$ choices) the $a_{i}$ are chosen as divisors of $\ell$ (at most $D(n)$ choices for each $i\in[k]$). ∎
@@ 3. The (1)-statement for two colours
For $p=\omega(n^{-1/9})$, the random set $n]_{p}$ a.a.s. contains a subset of the form (4).
@@ 4. More colours
This trick allows for colourings $\psi$ that avoid all triples of the forms $a+b=c$ and $a+b+1=c$ except $1+1=2$. Computer calculations show that with $3$ colours we can colour up to [^5] The construction here is AABBBACCCACCCABBBA. $18$ (while $S^{\prime}(3)=13$) and with four colours up to [^6] The construction here is AABBBACCCACCCABBBADDDADDDABBBADDDADDDABBBACCCACCCABBBA. $54$ (when $S^{\prime}(4)=40$). This improves the bounds for $r\in\{3,4\}$ to $n^{-1/19+o(1)}$ and $n^{-1/55+o(1)}$, respectively. However, we have not found a better general construction using this idea.
@@ 5. Concluding remarks
For the $r$-product-Schur problem with more colours, we showed that the threshold $p_{r}^{*}$ lies between $n^{-S^{\prime}(r)}$ and $n^{-S(r)^{2}}$ where $S(r)$ is the Schur number and $S^{\prime}(r)$ is the double-sum Schur number. We remark that one could also try to use the idea for our $2$-colour (0)-statement to get a lower bound in the $r$-colour case. Indeed, the idea there was to take a greedy colouring and show that it a.a.s. does not fail to colour all the integers. One can show by induction that there are $3^{r}$ integers in a (non-degenerate) configuration which witnesses the failure of the algorithm and so one cannot expect a better bound than $p_{r}^{*}\geq n^{-1/3^{r}}$ using this approach. As $S^{\prime}(r)\geq 3S(r-1)-1=\Omega(3.17176^{r})$ [AbbottHanson, SchurLowerBound], this does worse than the lower bound we give here for $r$ sufficiently large.
It would be very interesting to close the gap on our knowledge of $\log_{n}{1/p_{r}^{*}}$ which we know is between exponential and factorial in $r$. Given that this is also the limit of our knowledge on the growth of Schur numbers [SchurLowerBound, SchurUpperBound], this seems like it could be a considerable challenge. Nonetheless, it may be possible to improve the upper and lower bounds of $\log_{n}{1/p_{r}^{*}}$ in terms of the Schur number $S(r)$. Finally we remark that it would be natural to study other nonlinear equations in this random Ramsey setting and we believe our techniques developed here will be useful for this. In particular, some pertinent candidates are $a^{2}=bc$ and $ab=a+b$.
@@ Acknowledgements
Part of this work was done as part of the Barcelona Introduction to Mathematical Research – (BIMR 2024) and we would like to thank BGSMath and Centre de Recerca Matemàtica (CRM) for organising this summer program. We also thank Shagnik Das, Letícia Mattos and Juanjo Rué for interesting discussions around this topic.

### v2

@@ 2.1. Lack of forbidden configurations
For $n\in{\mathbb{N}}$, a function $p=p(n)\leq n^{-1/9}\omega(n)^{-1}$, and $S\sim[n]_{p}$ we have that a.a.s. there is not a forbidden configuration in $S$.
@@
For $n\in{\mathbb{N}}$, $p\leq n^{-1/9}\omega(n)^{-1}$, and $S\sim[n]_{p}$ we have that a.a.s. there is not a forbidden configuration in $S\cap(n^{1/9},n].$
@@
Let $e_{1},\dots,e_{k}\in\mathbb{N}$ with $1\leq e_{1}\leq e_{2}\leq\dots\leq e_{k}$ and $1\leq t\leq n$. Let $e\coloneq\sum_{i=1}^{k}e_{i}$. Then the number of $k$-tuples $(a_{1},\dots,a_{k})$ such that $a_{i}\in[n]$ for all $i\in[k]$, $a_{i}\geq t$ for all $i\in[k]$ with $e_{i}\geq 2$, and $\prod_{i=1}^{k}a_{i}^{e_{i}}\leq n$ is at most
@@
As $\prod_{i=1}^{k}a_{i}^{e_{i}}\leq n$, and $a_{i}\geq t$ for each $i\in[k]$ with $e_{i}\geq 2$, any valid choice of $(a_{1},\ldots,a_{k})$ will have $\ell:=\prod_{i=1}^{k}a_{i}\leq nt^{k-e}$. After choosing $\ell$ (at most $nt^{k-e}$ choices) the $a_{i}$ are chosen as divisors of $\ell$ (at most $D(n)$ choices for each $i\in[k]$). ∎
@@ 3. The (1)-statement for two colours
For $p=\omega(n^{-1/11})$, the random set $n]_{p}$ a.a.s. contains a subset of the form (4).
@@ 4. More colours
This trick allows for colourings $\psi$ that avoid all triples of the forms $a+b=c$ and $a+b+1=c$ except $1+1=2$. Computer calculations show that with $3$ colours we can colour up to [^5] The construction here is AABBBACCCACCCABBBA. $18$ (while $S^{\prime}(3)=13$) and with four colours up to [^6] The construction here is AABBBACCCACCCABBBADDDADDDABBBADDDADDDABBBACCCACCCABBBA. $54$ (when $S^{\prime}(4)=40$). This improves the bounds for $r\in\{3,4\}$ to $n^{-1/19}$ and $n^{-1/55}$, respectively. However, we have not found a better general construction using this idea.
@@ 5. Concluding remarks
For the $r$-product-Schur problem with more colours, we showed that the threshold $p_{r}^{*}$ lies between $n^{-S^{\prime}(r)}$ and $n^{-S(r)^{2}}$ where $S(r)$ is the Schur number and $S^{\prime}(r)$ is the double-sum Schur number. We remark that one could also try to use the idea for our $2$-colour (0)-statement to get a lower bound in the $r$-colour case. Indeed, the idea there was to take a greedy colouring and show that it a.a.s. does not fail to colour all the integers. One can show by induction that there are $3^{r}$ integers in a (non-degenerate) configuration which witnesses the failure of the algorithm and so one cannot expect a better bound than $p_{r}^{*}\geq n^{-1/3^{r}}$ using this approach. As $S^{\prime}(r)\geq 3S(r-1)-1=\Omega(3.17176^{r})$ [1, 11], this does worse than the lower bound we give here for $r$ sufficiently large.
It would be very interesting to close the gap on our knowledge of $\log_{n}{1/p_{r}^{*}}$ which we know is between exponential and factorial in $r$. Given that this is also the limit of our knowledge on the growth of Schur numbers [11, 30], this seems like it could be a considerable challenge. Nonetheless, it may be possible to improve the upper and lower bounds of $\log_{n}{1/p_{r}^{*}}$ in terms of the Schur number $S(r)$. Finally we remark that it would be natural to study other nonlinear equations in this random Ramsey setting and we believe our techniques developed here will be useful for this. In particular, some pertinent candidates are $a^{2}=bc$ (with $b\neq c$) and $ab=cd$ (with $\{a,b\}\neq\{c,d\}$).
@@ Acknowledgements
Part of this work was done during the Barcelona Introduction to Mathematical Research – (BIMR 2024) and we would like to thank BGSMath and Centre de Recerca Matemàtica (CRM) for organising this summer programme. We also thank Shagnik Das, Letícia Mattos and Juanjo Rué for interesting discussions around this topic, and Jonathan Chapman for helpful comments on a previous version of the manuscript.